OpenStudy (anonymous):

simply fraction in lowest terms. show work 56/119

jhonyy9 (jhonyy9):

8/17

OpenStudy (lollylau):

ah :)

OpenStudy (anonymous):

how did you get 8/17 explain

jhonyy9 (jhonyy9):

simplify by 7

jhonyy9 (jhonyy9):

8*7=56 17*7=119

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

so it would be 56 divided by 7 and 119 divided by 7

jhonyy9 (jhonyy9):

no will be 8*7 ---- = 8/17 17*7

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

okay thanks

jhonyy9 (jhonyy9):

good luck bye